SOLVE CUBIC EQUATION X^3 + 3X^2 - 28X = 0
Answer by venugopalramana(3286)

SOLVE CUBIC EQUATION X^3 + 3X^2 - 28X = 0
X(X^2+3X-28)=0
X[X^2+7X-4X-28]=0
X[X(X+7)-4(X+7)]=0
X(X-4)(X+7)=0
X=0.....OR......X=4.......OR.......X=-7